OpenStudy (anonymous):

25, 8, 10, 35, 5, 45, 40, 30, 20 What is the upper quartile of the data set above?A.35 B. 37.5 C. 37 D. 40

OpenStudy (anonymous):

IS it 40? @huntingman15 @Shawick @aum @yomamabf @undeadknight26 @marylou004

OpenStudy (anonymous):

I have no idea. Sorry :'(

OpenStudy (anonymous):

aw man i wanna see if i'm right ;w;

undeadknight26 (undeadknight26):

37.5

undeadknight26 (undeadknight26):

Using calculator: http://www.alcula.com/calculators/statistics/quartiles/

OpenStudy (anonymous):

ahhh thanks alot

OpenStudy (anonymous):

hows it 37.5 i got 40 somehow

undeadknight26 (undeadknight26):

Np my guests left so i said might as well :D Welcome to OS and please do stay :D

undeadknight26 (undeadknight26):

well first put them together in order. Can you do that for me mate?

OpenStudy (anonymous):

i don't think i can right atm i'm gonna pass out so i had to hurry n get help here

OpenStudy (anonymous):

im going to leave this up n nap

undeadknight26 (undeadknight26):

I wanna nap too D: Ill work it out den watch anime den nap :D

OpenStudy (anonymous):

ahh you should rest i'm a child with illnesses so i need it asap earlier i almost collasped in the shower

undeadknight26 (undeadknight26):

5, 8, 10, 20, 25, 30, 35, 40, 45.

undeadknight26 (undeadknight26):

Gur dun worry bout me XD Because you have illness gives me twice the reason to work and show you how to do it!

OpenStudy (anonymous):

you're a kind person thanks alot for working with me ^^ i'll explain the illness thing later but i can tell u now i'm legally blind in my left eye and need glases to see clearer i have tumors and joint pain problems all-day and night i have skin cancer and i get weak easily plus i'm 13 sorry i keep dozing off ;p

OpenStudy (anonymous):

do you have z a facebook?

undeadknight26 (undeadknight26):

median = 25 lower quartile = 9 (8 + 10 = 18/2 = 9) Upper quartile = 37.5 (30 + 45 = 75/2 = 37.5)

undeadknight26 (undeadknight26):

Nope! I have a G+ though! 5, 8, 10, 20, 25, 30, 35, 40, 45. median = 25 5, 8, 10, 20 8+10 = 18/2 = 9 30, 35, 40, 45 35+45 = 75/2 = 37.5

undeadknight26 (undeadknight26):

Boom boom n boom!